# Basic representations of genus zero nonabelian Hodge spaces

### v2

@@ 1.4. Main results: from enriched trees to classes of nearby representations
While we refer the reader to the body of the article for the detailed construction of the classes of representations, let us now briefly summarise their properties and discuss the link with the diagrams of [18]. Recall that if $(E,\nabla)$ is an algebraic connection on a Zariski open subset of $\mathbb{P}^{1}$, at each of its singularities it possesses a finite set of Stokes circles, which basically correspond to the different exponential factors $e^{q}$ appearing in the asymptotics of the horizontal sections of the connection close to the singularity. In our language they are encoded by the (modified) global irregular class $\bm{\breve{\Theta}}$ which is the collection of all Stokes circles of $(E,\nabla)$, with some integer multiplicities.
Theorem 1.6 (Prop. 3.14, 3.15).
[…]
@@
Let $(E,\nabla)$ be an irreducible irregular connection on $\mathbb{P}^{1}$ with enriched tree $(\mathbf{T},\bm{\mathcal{C}})$, and $(E^{\prime},\nabla^{\prime})\in\operatorname{SL}_{2}(\mathbb{C})\cdot(E,\nabla)$ an element in its orbit. We have:
@@
    There is a bijection between $N$ and the set of Stokes circles of $(E^{\prime},\nabla^{\prime})$. Furthermore:
@@
      If $(E^{\prime},\nabla^{\prime})$ is in the generic class of representations, all its Stokes circles are of generic form, i.e. are at infinity and are of slope $\leq 2$.
@@
      Otherwise, if $(E^{\prime},\nabla^{\prime})$ is in the $i$-th nongeneric class of representations for $i=1,\dots k$,
      • the Stokes circles corresponding to elements of $N_{i}^{-}$ via the bijection are at finite distance (i.e. different from $\infty\in\mathbb{P}^{1}$),
      • the Stokes circles corresponding to elements of $N_{i}^{+}$ are at infinity and of slope $>2$,
@@
[…]
@@
      $r_{i}=\sum_{I\in N_{i}^{+}}n_{I}R_{I}+\sum_{\begin{subarray}{c}l=1\\ l\neq i\end{subarray}}^{k}\left(\sum_{I\in N_{l}}n_{I}r_{I}\right).$
The link with the diagrams of [18] is as follows. Any irregular connection on $\mathbb{P}^{1}$ canonically determines both a diagram $\Gamma$ and an enriched tree $\mathscr{T}$, both invariant under symplectic transformations. Our main result says that the datum of the enriched tree allows us to construct $k+1$ different readings of the diagram, corresponding to the classes of representations in an $\operatorname{SL}_{2}(\mathbb{C})$-orbit. The enriched tree is thus a finer invariant that contains in general strictly more information than the diagram, i.e. we can view it as an enriched diagram. Concretely, the enriched tree determines explicitly the diagram, as well as a partition of the set of its core vertices into $k$-subsets.
@@
Let us also mention that the fission trees of [11] are closely related to combinatorial objects appearing in the study of the singularities of plane algebraic curves: they can be seen as a variant of the Eggers-Wall tree [46]. In some sense, we can thus view the short fission tree as a global analogue of Eggers-Wall trees, for the case of several singularities located at different points on $\mathbb{P}^{1}$.
@@
We apply our framework to the case of nonabelian Hodge spaces related to Painlevé equations, which have complex dimension 2 and constitute the simplest nontrivial examples of nontrivial nonabelian Hodge spaces. Concretely for all Painlevé moduli spaces we describe their classes of basic representations, which are classes of nearby representations of the standard rank 2 Lax representations known since the early 20th century [26, 27, 33]. In this way we recover many known alternative Lax representations as basic representations different from the standard one, and we can view the remaining basic representations as providing new Lax representations (in our abstract sense).
@@
While we refer the reader to the last section of the article for many more pictures of fission trees and basic representations for Painlevé moduli spaces, a concrete example is already summarized on Figure LABEL:fig:PIII_tree_and_diagram, which corresponds to the classes of basic representations of the Painlevé III moduli space which are the classes of nearby representations of the standard rank 2 Painlevé III Lax representation (see §4 for a more detailed discussion of this example). The figure illustrates how, starting from the short fission tree, the other classes of representations can be obtained, yielding different readings of the diagram.
@@
To conclude this introduction, let us sketch briefly sketch the main ideas leading to these results. An essential observation, already made in [7] in the simply-laced case, and in [18] in the general case, is that one can associate to any Stokes circle $I$ an element $\lambda(I)\in\mathbb{P}^{1}$ that we will call its Fourier sphere coefficient, such that the $\operatorname{SL}_{2}(\mathbb{C})$ action on the set of all possible Stokes circles (induced by the action on connections via the stationary phase formula) behaves in a very nice way with respect to those: it acts by homographies on the Fourier sphere $\mathbf{P}^{1}$, the copy of $\mathbb{P}^{1}$ where those coefficients live. Concretely, for
$A=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{SL}_{2}(\mathbb{C}),$
letting $h_{A}$ denote the corresponding homography
$h_{A}:z\mapsto\frac{az+b}{cz+d},$
we have that $A$ acts on the Fourier sphere coefficients by $h_{A}$.
@@
Using the results of [11] on admissible deformations, it is then quite straightforward to see that up to admissible deformations, the irregular class of $A\cdot(E,\nabla)$ only depends on whether we have $h_{A}(\lambda_{i})=\infty$ for some $i=1,\dots,k$, and if so for which $i$, which leads to the nearby representations: the generic representation corresponds to the case where $h_{A}(\lambda_{i})\neq\infty$ for all $i=1,\dots k$, and the $i$-th nongeneric representation to the case where $h_{A}(\lambda_{i})=\infty$. Each distinct Fourier sphere coefficient $\lambda_{i}$ corresponds to one principal subtree of the fission tree of the generic representation.
1.5. Outlook
@@ 1.6. Contents
Acknowledgements
I am funded by the PNRR Grant CF 44/14.11.2022, “Cohomological Hall Algebras of Smooth Surfaces and Applications”, led by Olivier Schiffmann. I also thank the funding of FCiências.ID, the great working environment of the Grupo de Física Matemática, where most of this work was carried out, and Philip Boalch, Gabriele Rembado, Giordano Cotti, Giulio Ruzza, Davide Masoero, Gabriele Degano, Julian Barragán Amado, for useful discussions.
2. Symplectic action on fission data of irregular connections on $\mathbb{P}^{1}$
@@
2.1. Formal data of irregular connections
Let us first recall the description of formal data of irregular connections on the Riemann sphere. The Turritin-Levelt theorem states that any connection on the formal punctured disc can be decomposed as a direct sum of elementary connections having only one exponential factor. This can be formulated in different ways; the one we will be using involves viewing exponential factors at a singularity as sections of a local system over the circle of directions around this singularity (see e.g. [18, 15] for more details).
2.1.1. The exponential local system
Let $\Sigma=\mathbb{P}^{1}$ and $a\in\Sigma$. Let $\pi_{a}:\widehat{\Sigma}_{a}\to\Sigma$ the real oriented blow-up at $a$ of $\Sigma$. The preimage $\partial_{a}:=\pi^{-1}(a)$ is a circle whose points correspond to the directions around $a$. An open subset of $\partial_{a}$ corresponds to a germ of sector at $a$.
Let $z_{a}$ be a local complex coordinate on $\Sigma$ in a neighbourhood of $a$, vanishing at $a$. The exponential local system $\mathcal{I}_{a}$ is a local system of sets (i.e. a covering space) over $\partial_{a}$ whose sections are germs of holomorphic functions on sectors of the form
$\sum_{i}b_{i}z_{a}^{-k_{i}},$
where $k_{i}\in\mathbb{Q}_{>0}$, and $b_{i}\in\mathbb{C}$. The connected component of such a local section is a finite order cover of the circle $\partial$. More precisely, let $r$ the smallest integer such that the expression $q=\sum_{i}b_{i}z_{a}^{-k_{i}}$ is a polynomial in $z_{a}^{-1/r}$. The corresponding holomorphic function is multivalued, and becomes single-valued when passing to a finite cover $t_{a}^{r}=z_{a}$. Therefore, the corresponding connected component, which we denote by $\langle q\rangle$, is a $r$-sheeted cover of $\partial_{a}$. As a topological space, it is homeomorphic to a circle, and $\mathcal{I}_{a}$ is thus a disjoint union of (an infinite number of) such Stokes circles.
The integer $r=:\operatorname{ram}(q)$ is the ramification order of the circle $I=\langle q\rangle$. If $r=1$ we say that the circle $\langle q\rangle$ is unramified.
If $d\in\partial_{a}$, we denote by $(\mathcal{I}_{a})_{d}$ the fibre of $\mathcal{I}$ over the direction $d$. Taking $d$ as a basis point of $\partial_{a}$, the monodromy of $\mathcal{I}_{a}$ is the automorphism
$\rho_{a}:(\mathcal{I}_{a})_{d}\to(\mathcal{I}_{a})_{d}$
of the fibre $(\mathcal{I}_{a})_{d}$ obtained when going once around $\partial_{a}$ in the positive direction.
Finally we define the global exponential local system as the disjoint union
$\mathcal{I}:=\sqcup_{a\in\Sigma}\mathcal{I}_{a}.$
It naturally comes with a projection $\pi:\mathcal{I}\to\partial$, where
$\partial:=\sqcup_{a\in\Sigma}\partial_{a}.$
@@ 2.1.1. The exponential local system
2.1.2. Local and global formal data
@@
An irregular class at $a$ is a function $\Theta_{a}:\pi_{0}(\mathcal{I}_{a})\to\mathbb{N}$ with compact support. The rank of an irregular class $\Theta_{a}$ is
@@
where $n_{I}=\Theta_{a}(I)$ is the multiplicity of the circle $I$. If $n_{I}\neq 0$ we say that $I$ is an active circle of $\Theta_{a}$, or a Stokes circle of $\Theta_{a}$.
The Turritin-Levelt theorem says that any rank $r$ connection on the formal punctured disc at $a$ uniquely determines a pair $(\Theta_{a},\mathcal{C}_{a})$, where $\Theta_{a}$ is a rank $r$ irregular class, and $\mathcal{C}_{a}$ is the datum of a conjugacy class $\mathcal{C}_{I}\in GL_{n_{I}}(\mathbb{C})$ for any active circle $I$ of $\Theta_{a}$ with multiplicity $n_{I}$. The pair $(\Theta_{a},\mathcal{C}_{a})$ constitutes the formal data of the connection and determines it completely up to isomorphism.
@@
The Stokes circles of $\Theta_{a}$ at each singularity $a\in\Sigma$ correspond to the exponential factors of the connection in the Turritin-Levelt normal form at $a$, and the conjugacy classes $\mathcal{C}_{a}$ encode its the formal monodromy. In particular, the connection has a regular singularity at $a$ if and only if its only Stokes circle at $a$ is the tame circle $\langle 0\rangle_{a}$.
@@
Notice that if we wanted to work with the full nonabelian Hodge package, we would need to consider weighted conjugacy classes (cf. [6]). Here for simplicity, since we won’t be trying in this work to prove that the different classes of representations correspond as one expects to isomorphisms between the full nonabelian Hodge spaces, we will just work will standard (unweighted) conjugacy classes.
2.1.3. Modified formal data
It turns out that the action of the Fourier transform on formal data of connections is more easily described in terms of modified formal data, obtained as follows: if $i$ is a tame circle at finite distance ($\pi(i)\neq\infty$ and $q_{i}=0$) then replace $n_{i}$ by the integer
$m_{i}=\text{rank}(A-1),$
where $A\in\mathcal{C}_{i}$. Thus $A=1+vu$ for a surjective map $u:\mathbb{C}^{n_{i}}\to\mathbb{C}^{m_{i}}$ and an injective map $v:\mathbb{C}^{m_{i}}\to\mathbb{C}^{n_{i}}$. Then replace $\mathcal{C}_{i}$ by the conjugacy class $\breve{\mathcal{C}}_{i}$ of $(1+uv)$ in $\operatorname{GL}_{m_{i}}(\mathbb{C})$; this class is called the child of $\mathcal{C}_{i}$ in [9, appendix]. If $\pi(i)=\infty$ or if $i$ is not tame ($q_{i}\neq 0$) then we do no modification: $m_{i}=n_{i}$ and $\breve{\mathcal{C}}_{i}=\mathcal{C}_{i}$. This defines modified formal data $(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}})$, where $\bm{\breve{\Theta}}$ is the collection of circles with multiplicities $m_{i}\geq 1$.

### v3

@@ 1.4 Main results: from enriched trees to classes of nearby representations
While we refer the reader to the body of the article for the detailed construction of the classes of representations (defined in Definition 3.10), let us now briefly summarise their properties and discuss the links with the diagrams of [20]. Recall that if $(E,\nabla)$ is an algebraic connection on a Zariski open subset of $\mathbb{P}^{1}$, at each of its singularities it possesses a finite set of Stokes circles, which basically correspond to the different exponential factors $e^{q}$ appearing in the asymptotics of the horizontal sections of the connection close to the singularity. In our language, they are encoded by the (modified) global irregular class $\bm{\breve{\Theta}}$ which is the collection of all Stokes circles of $(E,\nabla)$, with some integer multiplicities.
Theorem 1.6 (Propositions 3.15 and 3.16).
Keeping the setup of Theorem 1.5, for $i\in\{1,\dots,\allowbreak k\}$, let $\mathbf{T}_{i}$ be the $i$-th principal subtree of $\mathbf{T}$, and $N_{i}$ denote its set of leaves, so that we have a partition $N=N_{1}\sqcup\dots\sqcup N_{k}$. Furthermore, let $N_{i}^{+}\subset N_{i}$ be the subset of leaves in $N_{i}$ having an ancestor vertex in $\mathbf{T}_{i}$ of height $1<h<2$, and $N_{i}^{-}:=N_{i}\smallsetminus N_{i}^{+}$.
@@
Let $(E,\nabla)$ be an irreducible irregular connection on $\mathbb{P}^{1}$ with enriched tree $(\mathbf{T},\bm{\mathcal{C}})$, and $(E^{\prime},\nabla^{\prime})\allowbreak\in\operatorname{SL}_{2}(\mathbb{C})\cdot(E,\nabla)$ an element in its orbit. We have
@@ 1. $(1)$
    There is a bijection between $N$ and the set of Stokes circles of $(E^{\prime},\nabla^{\prime})$. Furthermore,
@@
      If $(E^{\prime},\nabla^{\prime})$ is in the generic class of representations, all its Stokes circles are of generic form, i.e., are at infinity and are of slope $\leq 2$.
@@
      Otherwise, if $(E^{\prime},\nabla^{\prime})$ is in the $i$-th nongeneric class of representations for $i\in\{1,\allowbreak\dots k\}$,
      • the Stokes circles corresponding to elements of $N_{i}^{-}$ via the bijection are at finite distance ($\big$(i.e., different from $\infty\in\mathbb{P}^{1}\big)$,
      • the Stokes circles corresponding to elements of $N_{i}^{+}$ are at infinity and of slope >2>2,
@@ 3. $(3)$
    The ranks of the classes of representations are as follows. For any leaf $I\in N_{i}$, let $r_{I}$ denote the lowest common multiple of the denominators ((when written in their reduced form)) of all the heights of its ancestors in $\mathbf{T}_{i}$, and let $n_{I}$ be the multiplicity of $I$. If $I\in N_{i}^{+}$, let $s_{I}$ be the integer such that $s_{I}/r_{I}$ is the greatest height $h$ of any ancestor of $I$ with $1<h<2$, and set $R_{I}:=s_{I}-r_{I}<r_{I}$.
@@
      $r_{i}=\sum_{I\in N_{i}^{+}}n_{I}R_{I}+\sum_{\begin{subarray}{c}l=1\\ l\neq i\end{subarray}}^{k}\Bigg(\sum_{I\in N_{l}}n_{I}r_{I}\Bigg).$
The link with the diagrams of [20] is as follows. Any irregular connection on $\mathbb{P}^{1}$ canonically determines both a diagram $\Gamma$ and an enriched tree $\mathscr{T}$, both invariant under symplectic transformations. Our main result says that the datum of the enriched tree allows us to construct k+1{k+1} different readings of the diagram, corresponding to the classes of representations in an $\operatorname{SL}_{2}(\mathbb{C})$-orbit. The enriched tree is thus a finer invariant that contains in general strictly more information than the diagram, i.e., we can view it as an enriched diagram. Concretely, the enriched tree determines explicitly the diagram, as well as a partition of the set of its core vertices into $k$ subsets.
@@
Let us also mention that the fission trees of [13] are closely related to combinatorial objects appearing in the study of the singularities of plane algebraic curves: they can be seen as a variant of the Eggers–Wall tree [51]. In some sense, we can thus view the short fission tree as a global analogue of Eggers–Wall trees, for the case of several singularities located at different points on $\mathbb{P}^{1}$.
@@
We apply our framework to the case of nonabelian Hodge spaces related to Painlevé equations, which have complex dimension 2 and constitute the simplest nontrivial examples of nonabelian Hodge spaces. Concretely, for all Painlevé moduli spaces we describe their classes of basic representations, which are classes of nearby representations of the standard rank 2 Lax representations known since the early 20th century [27, 29, 35]. In this way we recover many known alternative Lax representations as basic representations different from the standard ones, and we can view the remaining basic representations as providing new Lax representations (in our abstract sense).
@@
While we refer the reader to the last section of the article for many more pictures of fission trees and basic representations for Painlevé moduli spaces, a concrete example is already summarized on Figure 2, which corresponds to the classes of basic representations of the Painlevé III moduli space which are the classes of nearby representations of the standard rank 2 Painlevé III Lax representation (see Section 4.4 for a more detailed discussion of this example). The figure illustrates how, starting from the short fission tree, the other classes of representations can be obtained, yielding different readings of the diagram.
1st nongeneric rep.1/21/2rank 23 singularitiesKatz ranks 0, 0, 1/2Generic rep.11221/21/2rank 41 singularityKatz rank 2transform 1st subtreetransform 2nd subtree2nd nongeneric rep.1122rank 22 singularitiesKatz ranks 1, 1
Figure 2: An example: classes of basic representations of the Painlevé III moduli space, giving 3 readings of the same diagram. The rational numbers on the left of the fission forests are the heights of the vertices. The generic representation corresponds to a short fission tree with $k=2$ principal subtrees. Each leaf corresponds to a vertex in the diagram, the subset $N_{1}=N_{1}^{-}$ corresponds to blue vertices, and $N_{2}=N_{2}^{-}$ corresponds to the red vertex. There are thus 2 nongeneric classes of nearby representations, whose fission forests are drawn on the figure. For each $i\in\{1,2\}$ the forest is obtained from the short fission tree by changing the $i$-th principal subtree into its Fourier transform, and keeping the other principal subtree unchanged (up to some changes of truncations of the trees). Notice that the nongeneric representation on the right corresponds to the standard Painlevé III Lax representation, with two second order poles (Poincaré–Katz rank 1), while the one on the left corresponds to the alternative Lax representation for Painlevé sometimes referred to as ‘degenerate Painlevé V’, with one twisted irregular singularity (of Poincaré–Katz rank $1/2$), and two regular singularities. See Section 4.4, and in particular Remark 3.18, for further discussion of this example, including explicit expressions of irregular classes realizing these representations.
@@
To conclude this introduction, let us briefly sketch the main ideas leading to these results. An essential observation, already made in [8] in the simply-laced case, and in [20] in the general case, is that one can associate to any Stokes circle $I$ an element $\lambda(I)\in\mathbb{P}^{1}$ that we will call its Fourier sphere coefficient, such that the $\operatorname{SL}_{2}(\mathbb{C})$ action on the set of all possible Stokes circles (induced by the action on connections via the stationary phase formula) behaves in a very nice way with respect to those: it acts by homographies on the Fourier sphere $\mathbf{P}^{1}$, the copy of $\mathbb{P}^{1}$ where those coefficients live. Concretely, for $A=\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\in\operatorname{SL}_{2}(\mathbb{C})$, letting $h_{A}$ denote the corresponding homography $h_{A}\colon\ z\mapsto\frac{az+b}{cz+d}$, we have that $A$ acts on the Fourier sphere coefficients by $h_{A}$.
@@
Using the results of [13] on admissible deformations, it is then quite straightforward to see that up to admissible deformations, the irregular class of $A\cdot(E,\nabla)$ only depends on whether we have $h_{A}(\lambda_{i})=\infty$ for some $i\in\{1,\dots,k\}$, and if so for which $i$, which leads to the nearby representations: the generic class of representations corresponds to the case where $h_{A}(\lambda_{i})\neq\infty$ for all $i\in\{1,\dots k\}$ (hence the terminology “generic”), and the $i$-th nongeneric class of representations to the case where $h_{A}(\lambda_{i})=\infty$}. Each distinct Fourier sphere coefficient $\lambda_{i}$ corresponds to one principal subtree of the fission tree of the generic class of representations.
1.5 Outlook
@@ 1.6 Contents
2 Symplectic action on fission data of irregular connections on $\bm{\mathbb{P}^{1}}$
@@
2.1 Formal data of irregular connections
Let us first recall the description of formal data of irregular connections on the Riemann sphere.
The well-known Turritin–Levelt theorem states that any connection on the formal punctured disc can be decomposed as a direct sum of elementary connections having only one exponential factor. This can be formulated in different ways; the one we will be using is the point of view à la Deligne–Malgrange (see, e.g., [40]), where one views exponential factors at a singularity as sections of a local system over the circle of directions around this singularity (see, e.g., [12, 20] for more details).
2.1.1 The exponential local system
Let $\Sigma=\mathbb{P}^{1}$ and $a\in\Sigma$. Let $\pi_{a}\colon\widehat{\Sigma}_{a}\to\Sigma$ be the real oriented blow-up at $a$ of $\Sigma$. The preimage $\partial_{a}:=\pi^{-1}(a)$} is a circle whose points correspond to the directions around $a$. An open subset of $\partial_{a}$ corresponds to a germ of sector at $a$.
Let $z_{a}$ be a local complex coordinate on $\Sigma$ in a neighbourhood of $a$, vanishing at $a$. The exponential local system $\mathcal{I}_{a}$ is a local system of sets (i.e., a covering space) over $\partial_{a}$ whose sections are germs of holomorphic functions on sectors of the form $\sum_{i}b_{i}z_{a}^{-k_{i}}$, where $k_{i}\in\mathbb{Q}_{>0}$, and $b_{i}\in\mathbb{C}$. The connected component of such a local section is a finite order cover of the circle $\partial$. More precisely, let $r$ be the smallest integer such that the expression $q=\sum_{i}b_{i}z_{a}^{-k_{i}}$} is a polynomial in $z_{a}^{-1/r}$. The corresponding holomorphic function is multivalued, and becomes single-valued when passing to a finite cover $t_{a}^{r}=z_{a}$. Therefore, the corresponding connected component, which we denote by $\bigl\langle q\bigr\rangle_{a}$, is a $r$-sheeted cover of $\partial_{a}$. As a topological space, it is homeomorphic to a circle, and $\mathcal{I}_{a}$ is thus a disjoint union of (an infinite number of) such Stokes circles.
The integer $r$ is the ramification order of the Stokes circle $I=\bigl\langle q\bigr\rangle_{a}$, which we denote by $r=:\operatorname{ram}(q)=\operatorname{ram}(I)$. If $r=1$, we say that the Stokes circle $\bigl\langle q\bigr\rangle_{a}$ is untwisted, or unramified. The degree $s$ of $q$ as a polynomial is the irregularity $\operatorname{irr}(q)=\operatorname{irr}(I)$ of $q$ (or $I$), and the quotient $s/r$ is the slope of $q$ (or $I$), which we denote by $\operatorname{slope}(q)=\operatorname{slope}(I)$.
If $d\in\partial_{a}$, we denote by $(\mathcal{I}_{a})_{d}$ the fibre of $\mathcal{I}_{a}$ over the direction $d$. Taking $d$ as a basepoint of $\partial_{a}$, the monodromy of $\mathcal{I}_{a}$ is the automorphism $\rho_{a}\colon(\mathcal{I}_{a})_{d}\to(\mathcal{I}_{a})_{d}$ of the fibre $(\mathcal{I}_{a})_{d}$ obtained when going once around $\partial_{a}$ in the positive direction.
Finally, we define the global exponential local system as the disjoint union
$\mathcal{I}:=\bigsqcup_{a\in\Sigma}\mathcal{I}_{a}.$
It naturally comes with a projection $\pi\colon\mathcal{I}\to\partial$, where
$\partial:=\bigsqcup_{a\in\Sigma}\partial_{a}.$
@@ 2.1.1 The exponential local system
2.1.2 Local and global formal data
@@
A (local) irregular class at $a$ is a function $\Theta_{a}\colon\pi_{0}(\mathcal{I}_{a})\to\mathbb{N}$ with compact support. The rank of an irregular class $\Theta_{a}$ is
@@
where $n_{I}=\Theta_{a}(I)$ is the multiplicity of the circle $I$. If $n_{I}\neq 0$, we say that $I$ is an active circle of $\Theta_{a}$, or a Stokes circle of $\Theta_{a}$.
Furthermore, the Poincaré–Katz rank of $\Theta_{a}$ is
$\operatorname{Katz}(\Theta_{a}):=\max\left\{\operatorname{slope}(I)\mid n_{I}>0\right\}.$
The Turritin–Levelt theorem implies that any rank $r$ connection on the formal punctured disc $\operatorname{Spec}(\mathbb{C}((z_{a})))$ at $a$ uniquely determines a pair $(\Theta_{a},\mathcal{C}_{a})$, where $\Theta_{a}$ is a rank $r$ irregular class, and $\mathcal{C}_{a}$ is the datum of a conjugacy class $\mathcal{C}_{I}\in GL_{n_{I}}(\mathbb{C})$ for any active circle $I$ of $\Theta_{a}$ with multiplicity $n_{I}$. The pair $(\Theta_{a},\mathcal{C}_{a})$ constitutes the formal data of the connection and determines it completely up to isomorphism.
@@
The Stokes circles of $\Theta_{a}$ at each singularity $a\in\Sigma$ correspond to the exponential factors of the connection in the Turritin–Levelt normal form at $a$, and the conjugacy classes $\mathcal{C}_{a}$ encode its formal monodromy. In particular, the connection has a regular singularity at $a$ if and only if its only Stokes circle at $a$ is the tame circle $\bigl\langle 0\bigr\rangle_{a}$.
@@
Notice that if we wanted to work with the full nonabelian Hodge package, we would need to consider weighted conjugacy classes (cf. [7]). Here for simplicity, since we will not be trying in this work to prove that the different classes of representations correspond as one expects to isomorphisms between the full nonabelian Hodge spaces, we will just work with standard (unweighted) conjugacy classes.
2.1.3 Modified formal data
It turns out that the action of the Fourier transform on formal data of connections is more easily described in terms of modified formal data, obtained as follows: if $i$ is a tame circle at finite distance (that is $i=\bigl\langle 0\bigr\rangle_{a}$ with $a=\pi(i)\neq\infty$), then replace $n_{i}$ by the integer $m_{i}:=\text{rank}(A-1)$, where $A\in\mathcal{C}_{i}$. Thus $A=1+vu$ for a surjective map $u\colon\mathbb{C}^{n_{i}}\to\mathbb{C}^{m_{i}}$ and an injective map $v\colon\mathbb{C}^{m_{i}}\to\mathbb{C}^{n_{i}}$}. Then replace $\mathcal{C}_{i}$ by the conjugacy class $\breve{\mathcal{C}}_{i}$ of $(1+uv)$ in $\operatorname{GL}_{m_{i}}(\mathbb{C})$; this class is called the child of $\mathcal{C}_{i}$ in [10, Appendix]. If $\pi(i)=\infty$ or if $i$ is not tame ($i\neq\bigl\langle 0\bigr\rangle_{a}$ for some $a$), then we do no modification: $m_{i}:=n_{i}$ and $\breve{\mathcal{C}}_{i}:=\mathcal{C}_{i}$. This defines modified formal data $\bigl(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}}\bigr)$, where $\bm{\breve{\Theta}}$ is the collection of all Stokes circles with multiplicities $m_{i}\geq 1$.